A piece brass metal 170kg has it's temperature raised from 0degree to 30 degree.calculate it increase in volume green the density of brass at 0degree an 8.5 multiple by 10kgm raised to power -3 and it's cubic expansivity as 5.7 raised to power -5 multiple by 10 raised to power -5kelvin raised to power of-1
1
Expert's answer
2018-05-14T09:31:07-0400
ΔV=αVΔT
V=m/ρ
ΔV=α m/ρ ΔT=(5.7 ∙ 〖10〗^(-5) ) 170/((8.5∙ 〖10〗^3 ) ) (30-0)=3.4∙ 〖10〗^(-5) m^3.